### v1

@@ 2.3 Experimental Challenges
The experimental observation of MBS in Q1D systems presents significant difficulties. MBS can be detected by a variety of experimental methods, including the temperature dependence [39, 40, 23] and field dependence [16, 22] of the nuclear magnetic resonance (NMR) relaxation rate, field dependent wave vector measurements by neutron scattering experiments [12, 13, 14, 15, 35, 19] and field dependent magnetisation [16, 18]. Fig. 2d suggests that the MBS phase spans a wide range of magnetisation. However, the substantial quantum fluctuations in Q1D systems result in a very gradual increase in magnetisation at low fields. Thus, high magnetic fields are required to achieve sufficient magnetisation to stabilise MBS in these systems. Furthermore, while the binding energy of magnons in an isolated $J_{1}$-$J_{2}$ chain is already small, it decreases sharply with increasing AFM interchain interactions. As a result, the MBS is extremely sensitive to these AFM interchain interactions [24, 25]. Therefore, experimental setups often require conditions such as low temperatures and high magnetic fields to stabilise and observe MBS. Consequently, the practical observation of external field induced MBS remains a major challenge.
@@ 3 Magnon bound states by internal field
As previously discussed, the observation of MBS in Q1D systems has traditionally been a matter of contention, primarily due to the challenges in experimental detection. However, the novel mechanism proposed herein facilitates magnon binding via AFM interchain coupling, which can mimic the effect of an actual external magnetic field for each chain, thereby eliminating the need for an external magnetic field. This advancement presents a significant advantage for experimental observations. This section elucidates the underlying mechanism and demonstrates that MBS can indeed emerge within realistic parameter regimes.
@@ 3.1 Internal field from mean-field approximation of interchain couplings
In Q1D magnetic systems, the interplay between interchain and intrachain couplings plays a pivotal role in determining the magnetic ground state. Notably, when interchain couplings are significantly weaker than their intrachain counterparts, the interchain interactions can be effectively approximated through a mean-field approach. This approximation allows for the conceptualisation of the system as a collection of independent 1D chains influenced by an effective mean field. This concept was initially validated for isotropic, i.e., XXX, Heisenberg chains linked by perpendicular interchain couplings $J_{\perp}$, described by the Hamiltonian ${\cal H}=\sum_{i,j}\mbox{\boldmath$S$}_{i,j}\cdot\mbox{\boldmath$S$}_{i,j+1}+J_{\perp}\sum_{i,j}\mbox{\boldmath$S$}_{i,j}\cdot\mbox{\boldmath$S$}_{i+1,j}$ with $J_{\perp}>0$. In the case of $J_{\perp}=0$, the system lacks magnetic order and is classified as a Tomonaga-Luttinger liquid at zero temperature. However, the spin-spin correlation functions decay in a power law manner, indicating the presence of quasi-long-range order. Hence, if finite $J_{\perp}$ is switched on, the system exhibits Néel long-range order [41].
Assuming the order to be oriented along the $z$-direction in spin space, the Hamiltonian transforms into an effective single-chain problem described by ${\cal H}_{\rm eff}=J\sum_{i}\mbox{\boldmath$S$}_{i}\cdot\mbox{\boldmath$S$}_{i+1}-h_{\rm stag}\sum_{i}(-1)^{i}S^{z}_{i}+{\rm const.}$, where $h_{\rm stag}$ denotes the staggered field. This field, acting as an effective internal force, encapsulates the collective effect of the interactions with adjacent chains, inherent to the structure of the system. It elucidates the formation of Néel order in 2D square and 3D cubic lattices as a dimensional extension from a 1D chain. The configuration of this effective internal field depends on the structural arrangement of magnetic interactions within the system. For example, it can exhibit incommensurate oscillations if the net magnetisation of the system is nonzero [42].
@@
If the magnitude of the AFM interchain coupling is sufficiently large to induce FM ordering in each chain, the spins in adjacent chains will be polarised in opposite directions. This scenario represents a phase transition from a spiral state to a CAFO, governed by the AFM interchain coupling. By conceptualising the AFM interchain coupling as an effective internal field, each chain can be viewed as an FM-AFM $J_{1}$-$J_{2}$ chain under a uniform field, with the field direction being opposite in adjacent chains. If this mapping holds true, the effective Hamiltonian for a single chain should be expressible in a form similar to Equation (2). In other words, if the internal field induced by the AFM interchain coupling can break the spin symmetry of each chain, leading to a finite magnetisation, it might be possible to realise MBSs without the need for an external magnetic field. In the following subsections, we will numerically verify the formation of CAFO as the AFM interchain coupling is increased, and the realisation of MBSs.
3.2 Magnetisation in self-organized collinear antiferromagnetic ordering
Figure 3: Magnetization per site in the thermodynamic limit ($l_{x}\to\infty$) for various strengths of the XXZ anisotropy, evaluated at representative ratios $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$. The magnetization is presented as a function of the interchain couplings $J_{\rm a}$ and $J_{\rm b}$. Panels (a) and (b) illustrate the effects of introducing XXZ anisotropy on inchain and interchain couplings, respectively. The parameters used in the calculations of MES are marked with circles (see Fig. 4).
@@ 3.2 Magnetisation in self-organized collinear antiferromagnetic ordering
Given the bipartite lattice configuration perpendicular to the chains, we segregate the chain indices $j$ into even and odd numbers for each sublattice. For a comprehensive exploration of parameter dependencies, we confine our system to 2-chain cluster ($l_{y}=2$), a scenario that maximises quantum fluctuations, thereby posing a challenging environment for the establishment of the CAFO. Nonetheless, the formation of a CAFO under such stringent conditions suggests that more stable CAFO structures would naturally arise in systems with reduced quantum fluctuations, such as in two- or three-dimensional lattices. Indeed, we have confirmed a more pronounced CAFO stabilisation for 4-chain ($l_{y}=4$) and 8-chain ($l_{y}=8$) clusters illustrated in Fig. 1f in comparison to the $l_{y}=2$ case.
Using the DMRG method, we calculate the magnetisation per site in the thermodynamic limit ($l_{x}\to\infty$) for various strengths of the XXZ anisotropy and representative values of the ratio $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$. The results are plotted as a function of the interchain couplings $J_{\rm a}$ and $J_{\rm b}$ in Fig. 3. In Fig. 3a, the XXZ anisotropy ($\Delta_{1}<1$) is introduced in the intrachain coupling, while in Fig. 3b, the XXZ anisotropy ($\Delta_{\rm a,b}<1$) is included in the interchain coupling. In both cases, a critical interchain coupling exists for all combinations of $J_{2}/\lvert J_{1}\rvert$ and a kind of the interchain couplings, beyond which the system transitions to the CAFO phase. When $J_{1}$-$J_{2}$ chains in the spiral phase are coupled by frustrated AFM interchain couplings, increasing the interchain coupling reduces the propagation number. Once this number reaches zero, the system transitions to the CAFO phase [43]. Therefore, comparing the results for $J_{2}/\lvert J_{1}\rvert=1/3$ and $J_{2}/\lvert J_{1}\rvert=1/2$, the interchain coupling required to induce the transition to the CAFO phase is smaller for $J_{2}/\lvert J_{1}\rvert=1/3$, which originally has a smaller propagation number. Furthermore, as inferred from Eq. 4, the effect of $J_{\rm b}$ is generally larger than that of $J_{\rm a}$. For instance, for $J_{2}/\lvert J_{1}\rvert=1/3$, a slight XXZ anisotropy in the intrachain coupling leads to a significant magnetisation when the interchain coupling is $J_{\rm b}$. This scenario corresponds to the case of Li2CuO2 discussed in Sec. 4.1.
@@ 3.3 Magnon excitation spectrum
Figure 4: (a) Schematic illustration distinguishing a MBS from a non-MBS within the MES. (b,c) DDMRG results for the MES $G_{p}(\omega)$ at various values of $p=1$, $2$, $3$, and $4$, for representative ratios $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$, with the XXZ anisotropy applied to (b) inchain and (c) interchain couplings. (d) Angle-resolved MES $G_{1}(k,\omega)$ and $G_{2}(k,\omega)$ for $J_{2}/\lvert J_{1}\rvert=1/2$, $J_{\rm a}/\lvert J_{1}\rvert=1.0$, and $\Delta_{\rm a}=0.1$.
In real materials, the behaviour of coupled $J_{1}$-$J_{2}$ chains can be particularly fascinating when considering MBSs. A straightforward approach to detect these states is to calculate the binding energy of magnons, $E_{\rm b}(M,p)$, as done in Sec. 2.2 using Eq. (1). However, this approach faces a problem for coupled chains due to interchain exchange interactions, which lead to the non-conservation of the total $S^{z}$ within each chain. Consequently, Eq. (1) cannot be directly applied. To circumvent this issue, we investigate the presence of MBSs by calculating the magnon excitation spectrum (MES) for scenarios where $p$ spins are flipped simultaneously. The MES for $p$-magnon flips is defined as:
@@
We then investigate whether MBSs truly form when a CAFO is established by the internal field, as demonstrated in Sec. 3.2. In the CAFO, each $J_{1}$-$J_{2}$ chain orders ferromagnetically, with neighbouring $J_{1}$-$J_{2}$ chains having their magnetisations aligned anti-parallel along the $z$-axis. Here, we compute the MES for the $J_{1}$-$J_{2}$ chains with magnetisation pointing in the positive $z$-direction using Eq. (7). We note that an equivalent result is obtained for chains with magnetisation in the negative $z$-direction by replacing $S^{-}_{i}$ with $S^{+}_{i}$ in Eq. (7).
The MES computed using DDMRG method for various parameters is shown in Figs. 4b and 4c. The parameters used in those calculations are marked with circles in Fig. 3. To clarify the impact of CAFO on the MES, we compare the MES before and after the critical interchain coupling. In Fig. 4b, where $J_{1}$ is modified to include XXZ anisotropy, we observe that for $J_{2}/\lvert J_{1}\rvert=1/3$ and $\Delta_{1}=0.9$, no magnon binding is present at $J_{\rm b}=0.05$, as the lowest excitation is provided by $G_{1}(\omega)$. However, when $J_{\rm b}=0.1$, leading to the formation of CAFO, $G_{2}(\omega)$ provides the lowest excitation, indicating that the ground state is a nematic state. Intriguingly, the onset of $G_{3}(\omega)$ is very close to that of $G_{2}(\omega)$, suggesting that a triatic state exists very near the ground state. For $J_{2}/\lvert J_{1}\rvert=1/2$ and $\Delta_{1}=0.8$, the MES transitions from $G_{1}(\omega)$ to $G_{2}(\omega)$ as the lowest excitation when CAFO forms.
Fig. 4c shows the MES when XXZ anisotropy is introduced in the interchain coupling. Essentially, similar results are obtained as when XXZ anisotropy is applied to $J_{1}$. For $J_{2}/\lvert J_{1}\rvert=1/3$ and $\Delta_{\rm a}=0.1$, the ground state becomes a triatic state upon the formation of CAFO. For $J_{2}/\lvert J_{1}\rvert=1/2$ and $\Delta_{\rm a}=0.1$, the ground state becomes a nematic state upon CAFO. The value of $p$ for the ground state generally corresponds to the number of magnon bindings stabilised by an external field in a single $J_{1}$-$J_{2}$ chain for a given $J_{2}/\lvert J_{1}\rvert$. Given that $J_{2}/\lvert J_{1}\rvert=1/3$ is very close to the boundary between $p=2$ and $p=3$ at $J_{2}/\lvert J_{1}\rvert=0.367615$ in the high-field phase diagram of a single $J_{1}$-$J_{2}$ chain (see Figs. 2b, c, d), the MES shows $G_{2}(\omega)$ and $G_{3}(\omega)$ onsets close to each other, almost degenerate at the ground state. However, in any case, the presence of CAFO results in excitations significantly lower than those given by $G_{1}(\omega)$.
@@
$G_{p}(k,\omega)=\sum_{n>0}\lvert\langle n\lvert\sum_{j}\left(\prod_{i=j}^{j+p-1}S^{-}_{i}\right)\exp(ikr_{j})\lvert 0\rangle\lvert^{2}\delta(\omega-E_{n}+E_{0})$ (8)
@@ 4 Realisation of zero-field MBS in Q1D cuprates
Finally, we consider four materials as candidates for the potential formation of MBSs based on the mechanism proposed above.
@@ 4.1 Li2CuO2
The most promising material for the application of our MBS emergence mechanism is Li2CuO2. The magnetic couplings for this material are estimated to be $J_{1}=-228$ K, $J_{2}=75.7$ K, $J_{\rm a}=1$ K, and $J_{\rm b}=9$[44]. Also, paramagnetic and AFM resonance experiments have observed easy-axis anisotropy along the a-axis [45], which is estimated to be approximately $5\%$ of $J_{1}$ [46, 44, 47]. Therefore, we assume a small $J_{1}$ XXZ anisotropy with $\Delta_{1}=0.95$. Given that $J_{2}/\lvert J_{1}\rvert=0.332(>1/4)$, each $J_{1}$-$J_{2}$ chain would remain in the spiral phase if the interchain couplings were absent. However, the introduction of small interchain couplings $J_{\rm a}=1$ K and $J_{\rm b}=9$ K results in the system exhibiting a CAFO. It has been estimated that a $J_{\rm b}$ of $8.2$ K is required to induce CAFO [48], and the actual interchain coupling slightly exceeds this value. As shown in Fig. 1b, this system has a three-dimensional structure where each chain is coupled to four neighbouring chains through $J_{\rm a}$ and $J_{\rm b}$ (see Fig. 1d). Therefore, we perform our DMRG calculations using a $12\times 8$-chain cluster, applying open boundary conditions along the chain direction and periodic boundary conditions perpendicular to the chains, as depicted in Fig. 1f. Using the aforementioned parameters, we confirm a CAFO state with $2\lvert(\rvert\langle S^{z}_{i}\rangle)=0.965$. This magnitude closely matches the experimentally estimated Cu ion moment of $0.96(4)\mu_{\rm B}$ at $T=1.5$[49], further supported by ab-initio calculations [50, 51, 46].
We then examine the results of the MES calculations using DDMRG method. Fig. 5a plots $G_{p}(\omega)$ for $p$ values from 1 to 4. Focusing on a single $J_{1}$-$J_{2}$ chain, the FM order with a large magnetization of $0.9652\lvert\langle S^{z}_{i}\rangle\rvert=$ 0.965 resembles a state almost fully saturated by an external magnetic field. Consequently, one might expect an MBS to form. The spectra clearly show that the $p=3$ spectrum $G_{3}(\omega)$ has the lowest excitation at $\omega\sim 2.58$ meV, indicating a triatic ground state. The energy gap between this peak and the lowest energy peak of $G_{1}(\omega)$ at $\omega\sim 2.73$ meV is $\Delta_{\rm gap,1}\sim 0.15$ meV, corresponding to the magnon binding energy. Note that our DDMRG calculations are performed at $T=0$. Additionally, $G_{1}(\omega)$ exhibits another gap, $\Delta_{\rm gap,2}\sim 1.41$ meV, between $\omega\sim 2.73$ meV and $\omega\sim 4.17$ meV, likely attributed to the spin gap induced by XXZ anisotropy. However, since the ground state is not a simple singlet, $\Delta_{\rm gap,2}$ is not a perfect gap, and some in-gap states are present between $\omega\sim 2.73$ meV and $\omega\sim 4.17$ meV.

### v3

@@ 2.3 Experimental Challenges
The experimental observation of MBS in Q1D systems presents significant difficulties. MBS can be detected by a variety of experimental methods, including the temperature dependence [39, 40, 23] and field dependence [15, 22] of the nuclear magnetic resonance (NMR) relaxation rate, field dependent wave vector measurements by neutron scattering experiments [11, 12, 13, 14, 35, 19] and field dependent magnetisation [15, 18]. Fig. 2d suggests that the MBS phase spans a wide range of magnetisation. However, the substantial quantum fluctuations in Q1D systems result in a very gradual increase in magnetisation at low fields. Thus, high magnetic fields are required to achieve sufficient magnetisation to stabilize MBS in these systems. Furthermore, while the binding energy of magnons in an isolated $J_{1}$-$J_{2}$ chain is already small, it decreases sharply with increasing AFM interchain interactions. As a result, the MBS is extremely sensitive to these AFM interchain interactions [24, 25]. Therefore, experimental setups often require conditions such as low temperatures and high magnetic fields to stabilize and observe MBS. This results in keeping the practical observation of external field induced MBS a major challenge.
@@ 3 Magnon bound states by internal field
As previously discussed, the observation of MBS in Q1D systems has traditionally been a matter of contention, primarily due to the challenges in experimental detection. However, the novel mechanism proposed herein facilitates magnon binding via AFM interchain coupling, which can mimic the effect of an actual external magnetic field on each chain, thereby eliminating the need for an actual external magnetic field. This advancement presents a significant advantage for experimental observations. This section elucidates the underlying mechanism and demonstrates that MBS can indeed emerge within realistic parameter regimes.
@@ 3.1 Internal field from mean-field approximation of interchain couplings
In Q1D magnetic systems, the interplay between interchain and intrachain couplings plays a pivotal role in determining the magnetic ground state. Notably, when interchain couplings are significantly weaker than their intrachain counterparts, the interchain interactions can be effectively approximated through a mean-field approach. This approximation allows for the conceptualization of the system as a collection of independent 1D chains influenced by an effective mean field. This concept was initially validated for isotropic, i.e., XXX, Heisenberg chains linked by perpendicular interchain couplings $J_{\perp}$, described by the Hamiltonian ${\cal H}=\sum_{i,j}\mbox{\boldmath$S$}_{i,j}\cdot\mbox{\boldmath$S$}_{i,j+1}+J_{\perp}\sum_{i,j}\mbox{\boldmath$S$}_{i,j}\cdot\mbox{\boldmath$S$}_{i+1,j}$ with $J_{\perp}>0$. In the case of $J_{\perp}=0$, the system lacks magnetic order and is classified as a Tomonaga-Luttinger liquid at zero temperature. However, the spin-spin correlation functions decay in a power law manner, indicating the presence of quasi-long-range order. Hence, if finite $J_{\perp}$ is switched on, the system exhibits Néel long-range order [41].
Assuming the order to be oriented along the $z$-direction in spin space, the Hamiltonian transforms into an effective single-chain problem described by ${\cal H}_{\rm eff}=J\sum_{i}\mbox{\boldmath$S$}_{i}\cdot\mbox{\boldmath$S$}_{i+1}-h_{\rm stag}\sum_{i}(-1)^{i}S^{z}_{i}+{\rm const.}$, where $h_{\rm stag}$ denotes the staggered field. This field, acting as an effective internal force, encapsulates the collective effect of the interactions with adjacent chains, inherent to the structure of the system. It elucidates the formation of Néel order in 2D square and 3D cubic lattices as a dimensional extension from a 1D chain. The configuration of this effective internal field depends on the structural arrangement of magnetic interactions within the system. For example, it can exhibit incommensurate oscillations if the net magnetisation of the system is nonzero [42, 43].
@@
If the magnitude of the AFM interchain coupling is sufficiently large to induce FM ordering in each chain, the spins in adjacent chains will be polarised in opposite directions. This scenario represents a phase transition from a spiral state to a CAFO, governed by the AFM interchain coupling. By conceptualizing the AFM interchain coupling as an effective internal field, each chain can be viewed as an FM-AFM $J_{1}$-$J_{2}$ chain under a uniform field, with the field direction being opposite in adjacent chains. If this mapping holds true, the effective Hamiltonian for a single chain should be expressible in a form similar to Equation (2). In other words, if the internal field induced by the AFM interchain coupling can break the spin symmetry of each chain, leading to a finite magnetisation, it might be possible to realise MBSs without the need for an external magnetic field. In the following subsections, we will numerically verify the formation of CAFO as the AFM interchain coupling is increased, and the realisation of MBSs.
3.2 Magnetisation in self-organised collinear antiferromagnetic ordering
Figure 3: Magnetisation per site in the thermodynamic limit ($l_{x}\to\infty$) for various strengths of the XXZ anisotropy, evaluated at representative ratios $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$. The magnetisation is presented as a function of the interchain couplings $J_{\rm a}$ and $J_{\rm b}$. Panels (a) and (b) illustrate the effects of introducing XXZ anisotropy on inchain and interchain couplings, respectively. The parameters used in the calculations of MES are marked with circles (see Fig. 4).
@@ 3.2 Magnetisation in self-organised collinear antiferromagnetic ordering
Given the bipartite lattice configuration perpendicular to the chains, we segregate the chain indices $j$ into even and odd numbers for each sublattice. For a comprehensive exploration of parameter dependencies, we confine our system to 2-chain cluster ($l_{y}=2$), a scenario that maximizes quantum fluctuations, thereby posing a challenging environment for the establishment of the CAFO. The formation of a CAFO under such stringent conditions suggests that more stable CAFO structures would naturally arise in systems with reduced quantum fluctuations, such as in two- or three-dimensional lattices. Indeed, we have confirmed a more pronounced CAFO stabilisation for 4-chain ($l_{y}=4$) and 8-chain ($l_{y}=8$) clusters illustrated in Fig. 1f in comparison to the $l_{y}=2$ case.
Using the density-matrix renormalization group (DMRG) method, we calculate the magnetisation per site in the thermodynamic limit ($l_{x}\to\infty$) for various strengths of the XXZ anisotropy and representative values of the ratio $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$. Details on our DMRG calculations are given in Methods. The results are plotted as a function of the interchain couplings $J_{\rm a}$ and $J_{\rm b}$ in Fig. 3. In Fig. 3a, the XXZ anisotropy ($\Delta_{1}<1$) is introduced in the intrachain coupling, while in Fig. 3b, the XXZ anisotropy ($\Delta_{\rm a,b}<1$) is included in the interchain coupling. In both cases, a critical interchain coupling exists for all combinations of $J_{2}/\lvert J_{1}\rvert$ and a kind of the interchain couplings, beyond which the system transitions to the CAFO phase. When $J_{1}$-$J_{2}$ chains in the spiral phase are coupled by frustrated AFM interchain couplings, increasing the interchain coupling reduces the propagation number. Once this number reaches zero, the system transitions to the CAFO phase [44]. Therefore, comparing the results for $J_{2}/\lvert J_{1}\rvert=1/3$ and $J_{2}/\lvert J_{1}\rvert=1/2$, the interchain coupling required to induce the transition to the CAFO phase is smaller for $J_{2}/\lvert J_{1}\rvert=1/3$, which originally has a smaller propagation number. Furthermore, as inferred from Eq. 4, the effect of $J_{\rm b}$ is generally larger than that of $J_{\rm a}$. For instance, for $J_{2}/\lvert J_{1}\rvert=1/3$, a slight XXZ anisotropy in the intrachain coupling leads to a significant magnetisation when the interchain coupling is $J_{\rm b}$. This scenario corresponds to the case of Li2CuO2 discussed in Sec. 4.1.
@@ 3.3 Magnon excitation spectrum
Figure 4: (a) Schematic illustration distinguishing a MBS from a non-MBS within the MES framework. (b,c) DDMRG results for the MES $G_{p}(\omega)$ at various values of $p=1$, $2$, $3$, and $4$, for representative ratios $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$, with the XXZ anisotropy applied to (b) intrachain and (c) interchain couplings. (d) Angle-resolved MES $G_{1}(k,\omega)$ and $G_{2}(k,\omega)$ for $J_{2}/\lvert J_{1}\rvert=1/2$, $J_{\rm a}/\lvert J_{1}\rvert=1.0$, and $\Delta_{\rm a}=0.1$.
In real materials, the behavior of coupled $J_{1}$-$J_{2}$ chains can be particularly fascinating when considering MBSs. A straightforward approach to detect these states is to calculate the binding energy of magnons, $E_{\rm b}(M,p)$, as done in Sec. 2.2 using Eq. (1). However, this approach faces a problem for coupled chains due to interchain exchange interactions, which lead to the non-conservation of the total $S^{z}$ within each chain. Consequently, Eq. (1) cannot be directly applied. To circumvent this issue, we investigate the presence of MBSs by calculating the magnon excitation spectrum (MES) for scenarios where $p$ spins are flipped simultaneously. The MES for $p$-magnon flips is defined as:
@@
We then investigate whether MBSs truly form when a CAFO is established by the internal field, as demonstrated in Sec. 3.2. In the CAFO, each $J_{1}$-$J_{2}$ chain orders ferromagnetically, with neighboring $J_{1}$-$J_{2}$ chains having their magnetisations aligned anti-parallel along the $z$-axis. Here, we compute the MES for the $J_{1}$-$J_{2}$ chains with magnetisation pointing in the positive $z$-direction using Eq. (7). We note that an equivalent result is obtained for chains with magnetisation in the negative $z$-direction by replacing $S^{-}_{i}$ with $S^{+}_{i}$ in Eq. (7).
The MES computed using dynamical DMRG (DDMRG) method for various parameters is shown in Figs. 4b and 4c. Details on our DDMRG calculations are given in Methods. The parameters used in those calculations are marked with circles in Fig. 3. To clarify the impact of CAFO on the MES, we compare the MES before and after the critical interchain coupling. In Fig. 4b, where $J_{1}$ is modified to include XXZ anisotropy, we observe that for $J_{2}/\lvert J_{1}\rvert=1/3$ and $\Delta_{1}=0.9$, no magnon binding is present at $J_{\rm b}=0.05$, as the lowest excitation is provided by $G_{1}(\omega)$. However, when $J_{\rm b}=0.1$, leading to the formation of CAFO, $G_{2}(\omega)$ provides the lowest excitation, indicating that the ground state is a nematic state. Remarkably, the onset of $G_{3}(\omega)$ is very close to that of $G_{2}(\omega)$, suggesting that a triatic state exists very near the ground state. For $J_{2}/\lvert J_{1}\rvert=1/2$ and $\Delta_{1}=0.8$, the MES transitions from $G_{1}(\omega)$ to $G_{2}(\omega)$ as the lowest excitation when CAFO forms.
Fig. 4c shows the MES when XXZ anisotropy is introduced in the interchain coupling. Essentially, similar results are obtained as for XXZ anisotropy applied to $J_{1}$. For $J_{2}/\lvert J_{1}\rvert=1/3$ and $\Delta_{\rm a}=0.1$, the ground state becomes a triatic state upon the formation of CAFO. For $J_{2}/\lvert J_{1}\rvert=1/2$ and $\Delta_{\rm a}=0.1$, the ground state becomes a nematic state upon CAFO. The value of $p$ for the ground state generally corresponds to the number of magnon bindings stabilised by an external field in a single $J_{1}$-$J_{2}$ chain for a given $J_{2}/\lvert J_{1}\rvert$. Given that $J_{2}/\lvert J_{1}\rvert=1/3$ is very close to the boundary between $p=2$ and $p=3$ at $J_{2}/\lvert J_{1}\rvert=0.367615$ in the high-field phase diagram of a single $J_{1}$-$J_{2}$ chain (see Figs. 2b, c, d), the MES shows $G_{2}(\omega)$ and $G_{3}(\omega)$ onsets close to each other, almost degenerate at the ground state. However, in any case, the presence of CAFO results in excitations significantly lower than those given by $G_{1}(\omega)$.
@@
$G_{p}(k,\omega)=$ $\sum_{n>0}\lvert\langle n\lvert\sum_{j}\left(\prod_{i=j}^{j+p-1}S^{-}_{i}\right)\exp(ikr_{j})\lvert 0\rangle\lvert^{2}$
×$\times\delta(\omega-E_{n}+E_{0})$ (8)
@@ 4 Realisation of zero-field MBS in Q1D cuprates
Finally, we consider four materials candidates for the potential formation of MBSs based on the mechanism proposed above.
@@ 4.1 Li2CuO2
The most promising material for the application of our MBS emergence mechanism is Li2CuO2. The magnetic couplings for this material are estimated to be $J_{1}=-228$ K, $J_{2}=75.7$ K, $J_{\rm a}=1$ K, and $J_{\rm b}=9$[45]. Also, paramagnetic and AFM resonance experiments have observed easy-axis anisotropy along the a-axis [46], which is estimated to be approximately $5\%$ of $J_{1}$ [47, 45, 48]. Therefore, we assume a small $J_{1}$ XXZ anisotropy with $\Delta_{1}=0.95$. Given that $J_{2}/\lvert J_{1}\rvert=0.332(>1/4)$, each $J_{1}$-$J_{2}$ chain would remain in the spiral phase if the interchain couplings were absent. However, the introduction of small interchain couplings $J_{\rm a}=1$ K and $J_{\rm b}=9$ K results in the system exhibiting a CAFO. It has been estimated that a $J_{\rm b}$ of $8.2$ K is required to induce CAFO [49], and the actual interchain coupling slightly exceeds this value. As shown in Fig. 1b, this system has a three-dimensional structure where each chain is coupled to four neighboring chains through $J_{\rm a}$ and $J_{\rm b}$ (see Fig. 1d). Therefore, we perform our DMRG calculations using a $12\times 8$-chain cluster, applying open boundary conditions along the chain direction and periodic boundary conditions perpendicular to the chains, as depicted in Fig. 1f. Using the aforementioned parameters, we confirm a CAFO state with $0.9652\lvert(\langle S^{z}_{i}\rangle)\rvert=$ 0.965. This magnitude closely matches the experimentally estimated Cu ion moment of $0.96(4)\mu_{\rm B}$ at $T=1.5$[50], further supported by ab-initio calculations [51, 52, 47].
We then examine the results of the MES calculations using DDMRG method. Fig. 5a plots $G_{p}(\omega)$ for $p$ values from 1 to 4. Focusing on a single $J_{1}$-$J_{2}$ chain, the FM order with a large magnetisation of $0.9652\lvert\langle S^{z}_{i}\rangle\rvert=$ 0.965 resembles a state almost fully saturated by an external magnetic field. Consequently, one might expect an MBS to form. The spectra clearly show that the $p=3$ spectrum $G_{3}(\omega)$ has the lowest excitation at $\omega\sim 2.58$ meV, indicating a triatic ground state. The energy gap between this peak and the lowest energy peak of $G_{1}(\omega)$ at $\omega\sim 2.73$ meV is $\Delta_{\rm gap,1}\sim 0.15$ meV, corresponding to the magnon binding energy. Note that our DDMRG calculations are performed at $T=0$. Additionally, $G_{1}(\omega)$ exhibits another gap, $\Delta_{\rm gap,2}\sim 1.41$ meV, between $\omega\sim 2.73$ meV and $\omega\sim 4.17$ meV, likely attributed to the spin gap induced by XXZ anisotropy. However, since the ground state is not a simple singlet, $\Delta_{\rm gap,2}$ is not a perfect gap, and some in-gap states are present between $\omega\sim 2.73$ meV and $\omega\sim 4.17$ meV.